OpenStudy (anonymous):

Find the greatest common factor of 270 and 360.

OpenStudy (anonymous):

you try this answer 90

OpenStudy (anonymous):

I think it would b 270/2=135 360/2=180 idk

OpenStudy (anonymous):

270=3*3*3*3*2*5 360=3*3*2*2*2*5 the common factor therefore is:3*3*2*5=90